0.12gm of magnesium is treated with an acid which gives 0.60gm of anhydrous magnesium salt then equivalent weight of an acid is
1
Expert's answer
2019-07-02T04:38:53-0400

Solution.

n(Mg) = 0.005 mole

n(Mg) = n(salt) = 0.005 mole


"M(salt) = \frac{m}{n}"

M(salt) = 121.55


"M(acid) = M(salt)-M(Mg)"

M(acid) = 97.24


"M^{e} (acid) = \frac{97.24}{2}""M^e(acid) = 48.62"

Answer:

"M^e(acid) = 48.62"